OpenStudy (anonymous):

Solve for b in the formula 3 a + 2 b = c rewriting equations, INewton can you help?

OpenStudy (anonymous):

2b=c-3a b=(c-3a)/2 Is that what you want?

OpenStudy (anonymous):

yes. i think it is..... Thanks!

OpenStudy (anonymous):

you welcome. Become my fan :)

OpenStudy (anonymous):

be my fan!

OpenStudy (anonymous):

Thanks